write the equation of a line (in slope intercept form)/ that goes through (10,4) and has a slope of one-fifth
Answer by nerdybill(7384)

write the equation of a line (in slope intercept form)/ that goes through (10,4) and has a slope of one-fifth
.
plug given info into "point-slope" form:
y - y1 = m(x - x1)
y - 4 = (1/5)(x - 10)
y - 4 = (1/5)x - 2
y = (1/5)x + 2 (this is what they're looking for)